OpenStudy (anonymous):

Given the function g(x) = -3x + 4, compare and contrast g(-2) and g(4). Choose the statement that is true concerning these two values. Choices: The values of g(-2) and g(4) cannot be compared. The value of g(-2) is the same as the value of g(4). The value of g(-2) is larger than the value of g(4). The value of g(-2) is smaller than the value of g(4).

OpenStudy (anonymous):

First, find g(-2) and g(4). g(-2)=(-3)(-2) + 4 = 6+4 =10 g(4)=(-3)(4)+4 = -12+4 = -8 Then, compare the results. 10>-8 g(-2) > g(4) The value of g(-2) is larger than the value of g(4).